Calculate the after-tax cost of debt under each of the following conditions:

Interest rate of 13%; tax rate of 0%. Round your answer to two decimal places.

%

Interest rate of 13%; tax rate of 20%. Round your answer to two decimal places.

%

Interest rate of 13%; tax rate of 30%. Round your answer to two decimal places.

%
Answer #1

After-tax cost of debt = Cost of debt*(1-tax rate)

a) 13%*(1-0%)=13%

b) 13%*(1-20%) =10.40%

c) 13%*(1-30%) =9.10%